OpenStudy (anonymous):

First one to answer i will give you a medal :) I z+3 I<=7 A.)z<=-10 or z=>4 B.)-104 D.)-10<=z<=4

OpenStudy (anonymous):

20

OpenStudy (anonymous):

20 what?

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

I answered this already your answer is d

OpenStudy (anonymous):

d

OpenStudy (anonymous):

I z+3 I<=7 z+3<=7 subtract 3 from both sides z<=4

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

D

OpenStudy (anonymous):

With that you can see it eliminates answers A-C so your final answer is D

OpenStudy (anonymous):

wait what about this one?

OpenStudy (anonymous):

SO DO I GET A MEDAL

OpenStudy (anonymous):

you put 20 that wasnt a chocie

OpenStudy (anonymous):

what about what one?

OpenStudy (anonymous):

i said D.

OpenStudy (anonymous):

okay i will give you one lmao

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

I c-6 I<=6

OpenStudy (anonymous):

0<=C<=12

OpenStudy (anonymous):

A.)c<=0 or c=>12 B.)0<c<12 C.)0<=c<=12 D.)c<0 orc>12

OpenStudy (anonymous):

c

OpenStudy (anonymous):

You sure lol

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problem any time:)

OpenStudy (anonymous):

do you know how to tell if a equation is a function?

OpenStudy (anonymous):

Uhm idk I could try I kind of remember What does the problem look like